# Narrow line AGN selection in CEERS: spectroscopic selection, physical properties, X-ray and radio analysis

### v1

@@ Appendix C NLAGN Table
CEERS-613 | 6.7288 | C3He2 Ne5 N5 | 44.78 (H) | x | 8.28 | 3 | ||
CEERS-669 | 5.2714 | R3S2 | 45.5 ([OIII]) | ¿2.19 | 8.59 | 3 | ||
@@
CEERS-1400 | 4.4893 | M2 | x | x | 8.19 | 3 | ||
CEERS-1477 | 4.6307 | M1 M2 | 45.96 (H) | 1.3 | 8.54 | 3 | ||
CEERS-1536 | 5.0338 | M1 M2 | 45.45 (H) | 0.35 | 8.93 | 5 | ||
CEERS-1539 | 4.8838 | R3S2 R3O1 | 45.87 (H) | 1.3 | 9.03 | 10 | ||
CEERS-1561 | 6.1965 | M1 M2 | 45.47 (H) | 0.2 | 8.68 | 4 | ||
CEERS-1605 | 4.6305 | R3S2 M1 M2 | 45.14 (H) | 0.53 | 8.69 | 11 | ||
@@
CEERS-2668 | 2.8984 | R3N2 R3S2 | 45.9 ([OIII]) | ¿3.56 | 10.42 | 19 | ||
CEERS-2754 | 2.238 | He2N2 | x | 1.58 | x | x | ||
CEERS-2808 | 3.3834 | R3S2 | 45.18 (H) | 1.5 | 10.68 | 619 | X-ray | |
CEERS-2900 | 1.9004 | R3N2 R3S2 | 43.88 ([OIII]) | ¿2.21 | 11.17 | 0 | X-ray & Radio | |
CEERS-2904 | 2.0249 | R3N2 R3S2 R3O1 He2N2 M1 M2 | 45.13 (H) | ¿2.33 | 10.78 | 26 | X-ray & BLAGN | |
CEERS-2919 | 1.7177 | R3N2 R3S2 R3O1 He2N2 | 45.02 (H) | 1.88 | 10.62 | 203 | X-ray & BLAGN | |
CEERS-3129 | 1.0373 | R3N2 R3S2 He2N2 | 44.68 (H) | 2.28 | 11.04 | 309 | Radio & BLAGN | |
CEERS-3187 | 1.4896 | He2N2 | 43.57 (H) | -0.34 | 8.23 | 3 | ||
CEERS-3223 | 2.2862 | R3S2 | 43.93 (H) | 0.19 | 8.9 | 1 | ||
CEERS-3535 | 2.4886 | R3S2 | 44.97 (H) | 0.15 | 8.87 | 6 | ||
CEERS-3585 | 3.8662 | M2 | 46.01 (H) | 2.02 | 7.97 | 2 | ||
CEERS-3838 | 1.9221 | R3S2 R3O1 | 44.9 (H) | 1.94 | 9.16 | 11 | ||
CEERS-4210 | 5.2549 | Ne4 | 45.53 (H) | 0.64 | 8.69 | 12 | ||
CEERS-4385 | 3.4219 | He2N2 | 45.28 (H) | 0.74 | 9.33 | 15 | ||
CEERS-4402 | 3.3293 | R3O1 | 45.28 (H) | 0.51 | 9.82 | 23 | ||
CEERS-5117 | 1.5728 | R3S2 R3O1 | 44.56 (H) | 1.24 | 9.26 | 2 | ||
CEERS-6689 | 1.4608 | He2N2 | x | ¿5.94 | 10.82 | 95 | ||
CEERS-8041 | 2.5319 | R3S2 | 45.3 ([OIII]) | ¿2.93 | 9.24 | 8 | ||
CEERS-8110 | 1.5662 | R3O1 | 44.48 (H) | 0.46 | 9.72 | 5 | ||
CEERS-8299 | 2.1567 | Ne5 | 45.91 (H) | 2.54 | 10.1 | 5 | ||
CEERS-9156 | 2.0534 | R3S2 | 45.1 (H) | 0.45 | 9.54 | 6 | ||
CEERS-9238 | 1.7415 | R3S2 | 45.28 (H) | 1.8 | 9.29 | 3 | ||
CEERS-13661 | 3.3279 | R3S2 Ne4 | 45.11 (H) | 0.88 | 9.03 | 9 | ||
CEERS-16551 | 2.5714 | R3S2 R3O1 | 44.35 (H) | -0.46 | 8.57 | 5 | ||
CEERS-17496 | 2.3732 | R3S2 R3O1 | 44.46 (H) | -0.29 | 8.79 | 3 | ||
CEERS-31075 | 1.2831 | He2N2 | 44.42 (H) | 0.12 | 7.8 | 0 | ||

### v3

@@ Appendix C Final narrow-line AGN table
| CEERS-613* | 6.7288 | C3He2 | 44.78 (H) | x | 8.28 | 3 | ||
| CEERS-669 | 5.2714 | R3S2 | 45.22 ([OIII]) | ¿1.92 | 8.59 | 3 | ||
@@
| CEERS-1400 | 4.4893 | M2 | x | 0 | 8.19 | 3 | ||
| CEERS-1477 | 4.6307 | M1 M2 | 45.8 (H) | 1.14 | 8.54 | 3 | ||
| CEERS-1536 | 5.0338 | M1 M2 | 45.41 (H) | 0.3 | 8.93 | 5 | ||
| CEERS-1539 | 4.8838 | R3S2 R3O1 | 45.71 (H) | 1.14 | 9.03 | 10 | ||
| CEERS-1561 | 6.1965 | M1 M2 | 45.45 (H) | 0.18 | 8.68 | 4 | ||
| CEERS-1605 | 4.6305 | R3S2 M1 M2 | 45.07 (H) | 0.46 | 8.69 | 11 | ||
@@
| CEERS-2668* | 2.8984 | R3N2 R3S2 | 45.5 ([OIII]) | ¿3.32 | 10.42 | 19 | ||
| CEERS-2754 | 2.238 | He2N2 | x | 1.47 | x | x | ||
| CEERS-2808 | 3.3834 | R3S2 | 44.99 (H) | 1.31 | 10.68 | 619 | X-ray | |
| CEERS-2900 | 1.9004 | R3N2 R3S2 | 43.62 ([OIII]) | ¿2.06 | 11.17 | 0 | X-ray & Radio | |
| CEERS-2904 | 2.0249 | R3N2 R3S2 R3O1 He2N2 M1 M2 | 44.9 (H) | ¿2.17 | 10.78 | 26 | X-ray & BLAGN | |
| CEERS-2919 | 1.7177 | R3N2 R3S2 R3O1 He2N2 | 44.84 (H) | 1.75 | 10.62 | 203 | X-ray & BLAGN | |
| CEERS-3129 | 1.0373 | R3N2 R3S2 He2N2 | 44.46 (H) | 2.12 | 11.04 | 309 | Radio & BLAGN | |
| CEERS-3187 | 1.4896 | He2N2 | 43.76 (H) | 0 | 8.23 | 3 | ||
| CEERS-3223 | 2.2862 | R3S2 | 43.91 (H) | 0.18 | 8.9 | 1 | ||
| CEERS-3535* | 2.4886 | R3S2 | 44.95 (H) | 0.14 | 8.87 | 6 | ||
| CEERS-3585* | 3.8662 | M2 | 45.76 (H) | 1.77 | 7.97 | 2 | ||
| CEERS-3838 | 1.9221 | R3S2 R3O1 | 44.68 (H) | 1.8 | 9.16 | 11 | ||
| CEERS-4210 | 5.2549 | Ne4 | 45.45 (H) | 0.56 | 8.69 | 12 | ||
| CEERS-4385 | 3.4219 | He2N2 | 45.19 (H) | 0.65 | 9.33 | 15 | ||
| CEERS-4402 | 3.3293 | R3O1 | 45.21 (H) | 0.44 | 9.82 | 23 | ||
| CEERS-5117 | 1.5728 | R3S2 R3O1 | 44.41 (H) | 1.16 | 9.26 | 2 | ||
| CEERS-6689 | 1.4608 | He2N2 | x | ¿5.53 | 10.82 | 95 | ||
| CEERS-8041 | 2.5319 | R3S2 | 44.96 ([OIII]) | ¿2.73 | 9.24 | 8 | ||
| CEERS-8110 | 1.5662 | R3O1 | 44.43 (H) | 0.43 | 9.72 | 5 | ||
| CEERS-8299 | 2.1567 | Ne5 | 45.62 (H) | 2.37 | 10.1 | 5 | ||
| CEERS-9156 | 2.0534 | R3S2 | 45.05 (H) | 0.42 | 9.54 | 6 | ||
| CEERS-9238 | 1.7415 | R3S2 | 45.07 (H) | 1.67 | 9.29 | 3 | ||
| CEERS-13661 | 3.3279 | R3S2 Ne4 | 45.0 (H) | 0.76 | 9.03 | 9 | ||
| CEERS-16551 | 2.5714 | R3S2 R3O1 | 44.61 (H) | 0 | 8.57 | 5 | ||
| CEERS-17496 | 2.3732 | R3S2 R3O1 | 44.62 (H) | 0 | 8.79 | 3 | ||
| CEERS-31075 | 1.2831 | He2N2 | 44.4 (H) | 0.11 | 7.8 | 0 | ||